the frequency of revolution of electron in nth bohr orbit is Vn. Then what is the graph between log n and log (Vn/V1) ?

Grade:12th pass

1 Answers

Mayank
5 years ago
Vn is directly proportional to Z2/n3 then taking Z=1 , we have Vn is inversely proprotional to n  so Vn = k/n ^3 where k is a constant so putting n=1 we have V1 =k now log(Vn/V1)=log(1/n^3) =-3log(n) so graph of log(Vn/V1) and log (n) is a straight line passing through origin with slope -3